OpenStudy (anonymous):

Rita owes $9,739 on a credit card with a 21.5% interest rate compounded monthly. What monthly payment should she make to pay off this debt in six years, assuming she does not charge any more purchases with the card? $909.17 $135.26 $241.82 $2,093.89

OpenStudy (dumbcow):

its 241.82 is this a finance class? or a math class?

OpenStudy (anonymous):

Finance, how did you get that answer?

OpenStudy (dumbcow):

i used a finance calculator also i notice these answer choices are not even close you could get right answer by elimination

OpenStudy (dumbcow):

assuming no interest....9739/72 = 135.26 obviously monthly payment will be slightly larger than this

OpenStudy (dumbcow):

if you want to use a formula, memorize this: \[P = B \frac{1-v}{v(1-v^{n})}\] where \[v = \frac{1}{1+i/12}\]